The cell potential of a redox reaction occurring in an electrochemical cell under any set of temperature and concentration conditions can be determined from the standard cell potential of the cell using the Nernst equation where E is the cell potential of the cell, E° is the standard cell potential of the cell, R is the gas constant, T is the temperature in kelvin, n is the moles of electrons transferred in the reaction, and Q is the reaction quotient. Use this relationship to answer the problem below.

For the following oxidation-reduction reaction 3 Cu2+(aq) + 2 Au(s) → 3 Cu(s) + 2 Au3+(aq) the standard cell potential is −1.16 V. What is the actual cell potential of the cell if the temperature is 318 K, the initial Cu2+ concentration is 0.00113 M, and the initial Au3+ concentration is 0.000145 M? (Note that the reaction involves the transfer of 6 moles of electrons, and the reaction quotient is 14.6.)
Answer #1

Ecell = -1-16V 34 2+ + 2 All 3 cut 2 A43t E = Cell - 08.591 log lp = -1.16 -0.0591 Log (Aust] [unt 273 = -1-16-0059) log (6.0